OpenStudy (anonymous):

the area of the trapezoid is A=0.5(b1+b2)h find b1 if A=100 h=5 b2=12

OpenStudy (campbell_st):

substitute your information into the formula \[100 = 0.5\times(b _{1} + 12)\times12..... 100 = 6\times(b _{1} + 12)\] now solve for x the first step is to divide both sides by 6 \[\frac{100}{6} = b _{1} + 12\] I'll leave you to finish the calculations

OpenStudy (campbell_st):

oops should be b1 not x

OpenStudy (anonymous):

kay thanx [: